The following data has been taken from the accounting records of Chase Corporation for the year ended 2019:

Sales                                                                $930,000

Raw materials inventory, beginning                    $70,000                       

Raw materials inventory, ending             $40,000

Raw materials, purchases                               $190,000

Direct labor                                                     $150,000

Manufacturing overhead                                 $210,000

Administrative expenses                                $95,000

Selling expenses                                             $120,000

Work in process inventory, beginning        $80,000

Work in process inventory, ending             $75,000

Finished goods inventory, beginning                    $90,000

Finished goods inventory, ending             $140,000

Use the data to answer the following questions (a - c)

    1. What was the cost of the raw materials used in production during the year?
    1. What was the cost of goods manufactured (finished) for the year?
    1. What was the cost of goods sold for the year?
Answer #1

a) Cost of raw material used = 70000+190000-40000 = 220000

b) Cost of goods manufactured = 220000+150000+210000+80000-75000 = $585000

c) Cost of goods sold = 90000+585000-140000 = 535000